OpenStudy (calculusxy):

Susan begins counting backward from 1298 by 4's, saying one number every 5 seconds. At the same time, Jim begins counting forward from 171 by 3's, saying one number every 5 seconds. What number will they both say at the same time? @dumbcow @Boblovesmath

OpenStudy (calculusxy):

@hartnn

OpenStudy (anonymous):

you should equate them: 1298 - 4(x) = 171 +3x and then you solve for x and you plug it back into the equation to get the number you want. The reason why we equate them is that the number they lay upon should be the same.

OpenStudy (calculusxy):

@Boblovesmath Can you help me equate this equation by solving for x. @hartnn Can you also help me if possible, :)

OpenStudy (anonymous):

I just equated it for you it was the first thing that I wrote.

OpenStudy (calculusxy):

But what is the number? There is a specific number to this problem.

OpenStudy (anonymous):

yeah you find x and then to find the number you just plug it into to either 1298 - 4x or 171 + 3x as you will get the same number. All the best.

OpenStudy (calculusxy):

@hartnn Please help

hartnn (hartnn):

yes, i think 1298 - 4(x) = 171 +3x would work, because the time interval for both (5 sec) is same

OpenStudy (calculusxy):

Should I give you the answer choices?

hartnn (hartnn):

is 654 in it ?

OpenStudy (calculusxy):

Yes. How how did u get that? Please help me.

hartnn (hartnn):

did u get how that equation is formed ?

OpenStudy (calculusxy):

Kind of. Can u explain the usage for x in it.

hartnn (hartnn):

after 5 sec, you have decremented 1298 by 4, after 10 sec , decremented 1298 by 8 and so on, after x seconds, which we consider the time when we have same count, you would have decremented 1298 by 4x which is 1298 -4x

OpenStudy (calculusxy):

Ok now how did u get 654 from that equation?

hartnn (hartnn):

1298-4x = 171+3x add 4x on both sides then subtract 171 from both sides

OpenStudy (calculusxy):

So do I add 4x to 3x?

hartnn (hartnn):

yes

OpenStudy (calculusxy):

So that would be 1298=171+7x?

hartnn (hartnn):

yes

OpenStudy (calculusxy):

Then I do 1298-171, the difference divided by 7 will equal to x. Right?

hartnn (hartnn):

yes

hartnn (hartnn):

but the number you require is 1298 - 4x or 171 +3x

OpenStudy (calculusxy):

But wouldn't the division be equal to 654?

hartnn (hartnn):

which division ?

OpenStudy (calculusxy):

Difference divided by 4?

hartnn (hartnn):

no...it won't did u do it ?

OpenStudy (calculusxy):

*Sorry by 7 :/

OpenStudy (calculusxy):

I don't have a calculator at hand.

hartnn (hartnn):

1298 - 171 = 1127 1127/7 = 161

hartnn (hartnn):

171 + 3*161 = 654

OpenStudy (calculusxy):

So what do I do now?..??

hartnn (hartnn):

where are you even stuck ?

OpenStudy (calculusxy):

Ugh! By why 3??

hartnn (hartnn):

171 +3x is your number

OpenStudy (calculusxy):

Thx..

hartnn (hartnn):

welcome ^_^